OpenStudy (anonymous):

how do you write thirty six squared? And does 36² mean 36 x 36 or does it mean 6 x 6?

OpenStudy (anonymous):

36x36

geerky42 (geerky42):

How do you write thirty six squared? 36² And does 36² mean 36 x 36 or does it mean 6 x 6? When you square something, you multiply this number by itself, so \(36^2=36\times36\)

sammixboo (sammixboo):

Mmhmm \(\sf \large \color{magenta}{36^2~=36~x~36}\) \(\sf \large \color{magenta}{36^3~=36~x~36~x~36}\) \(\sf \large \color{magenta}{36^4~=36~x~36~x~36~x~36}\) \(\sf \large \color{magenta}{36^5~=36~x~36~x~36~x~36~x~36}\) \(\sf \large \color{magenta}{36^6~=36~x~36~x~36~x~36~x~36~x~36}\) And so on

sammixboo (sammixboo):

And \(\sf \large \color{magenta}{36^2~=~Squared}\) \(\sf \large \color{magenta}{36^3~=~Cubed}\)